OpenStudy (anonymous):

Write the system of inequalities that accompanies your graph the graph in 10 by 10

OpenStudy (inkyvoyd):

You do realize that without the system of inequalities we can't even begin to help you right?

OpenStudy (anonymous):

I need help makin the system of inequalities i have no idea how to make them

OpenStudy (inkyvoyd):

http://www.purplemath.com/modules/syslneq.htm Tell me if this helps.

OpenStudy (anonymous):

I just need to know how to make a system of inequalities for a 10 x 10 graph i am really confused now :\

OpenStudy (inkyvoyd):

Read it, it helps.

OpenStudy (inkyvoyd):

Hey, did you read the link?

OpenStudy (anonymous):

I read it and i am more confused :( i just need to know how to make a inequalitie for my graph ....................

OpenStudy (amistre64):

inequalities define an area bordered by a line or curve.

OpenStudy (amistre64):

write an equation: then change the = sign to either > or <

OpenStudy (anonymous):

how do i writ a equasion? sorry guys i know nothing about this subject :(

OpenStudy (amistre64):

usually by making up something to put on the left side of an equals sign; and something to equate it to on the right side of an equals sign: y = 2x + 1 is an equation z = cos(x) + tan(x+y) is an equation etc ....

OpenStudy (anonymous):

so if my graph is 10 by 10 my equation can be y=-4+3? that is just one

OpenStudy (amistre64):

it can be yes, or simplify that to y = -1

OpenStudy (amistre64):

then change the equal sign to either > or < to make it an inequality

OpenStudy (anonymous):

!!! ok!!! thank u now i know what to do!

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

one more question if this is my inqualtie y>4+1 and this is my graph how would i shade it?